# UDKAG: Augmenting Large Vision-Language Models with Up-to-Date Knowledge

### v1

@@ 5.3 Ablation Studies
The results of ablation studies on UDK-VQA are shown in Table 2. Firstly, we investigate simple IAG methods, including using the similarity between questions and segments to select segments, i.e., IAG (SIM $Q$), using the similarity between images and segments to select segments i.e., IAG (SIM $V$), using the averaged similarity of the the above two similarities to select segments, i.e., IAG (SIM $QV$). These methods show limited improvements and achieve unsatisfactory accuracy.
@@ 5.6 Analysis of Website Filter
6 Limitations
In our implementation, the hierarchical filtering model is trained apart from the LVLMs, which may affect the performance considering the differences in training data, architecture design, and abilities of LVLMs. In the future, we will consider training the filter and the LVLMs in an end-to-end manner.
@@ 7 Conclusion
In this work, we have presented UDKAG, a plug-and-play framework to augment LVLMs in handling visual question answering about up-to-date knowledge. By introducing a hierarchical filtering model, the framework enables LVLMs to access up-to-date knowledge. A UDK-VQA dataset is further curated by scraping up-to-date news and generating news-related VQA samples. The dataset enables quantitatively evaluate the ability of LVLMs to respond to questions about up-to-date knowledge. Experimental results on UDK-VQA demonstrate that our framework can significantly boost the performance of LVLMs for answering questions requiring up-to-date knowledge.

### v2

@@ 5.3 Ablation Studies
The experimental results of ablation studies on the proposed UDK-VQA dataset are shown in Table 3, where we use LLaVA-1.6 [23] as the baseline. Firstly, we investigate simple IAG methods, including using the similarity between questions and segments to select segments, i.e., IAG (SIM $Q$), using the similarity between images and segments to select segments i.e., IAG (SIM $V$), using the averaged similarity of the the above two similarities to select segments, i.e., IAG (SIM $QV$). These methods show limited improvements and achieve unsatisfactory accuracy.
@@ 5.4 Analysis of Pseudo-Score Generation
5.5 Analysis of Training Strategy
Table 4: Experiments of different training strategies on UDK-VQA.
LVLM Variant Training Acc
Strategy (%)
QWen-VL Raw - 35.2
Ours Joint 68.5
Ours Separate 84.8
LLaVA Raw - 41.2
1.5 Ours Joint 68.0
Ours Separate 88.9
Would jointly training the hierarchical filtering model with LVLMs result in greater improvements? We use LLaVA-1.5 [37] as the backbone for the hierarchical filtering model and conduct experiments with Qwen-VL [27] and LLaVA-1.5 as the LVLMs. As shown Table 4, separate training, where the LVLMs are frozen during the training of the hierarchical filtering model, leads to a more significant improvement in performance. The main reasons are: (1) Our training data uses pseudo-labeling instead of high-quality human annotation. Training based on such data may cause LVLMs to lose their original semantic understanding capabilities. (2) Our training and testing sets are generated from news from different time periods, involving different entities and having different distributions. Training LVLMs on our training set easily leads to overfitting, resulting in lower generalization on the test set.
@@ 5.7 Analysis of Website Filter
5.8 Analysis of Snippet Completeness
The incomplete snippets returned by search engines lead us to consider whether providing complete snippets to the website filter could result in further improvements. We present the experimental results of snippet completeness on our UDK-VQA dataset in the Table 5, where $\theta$ represents the percentage as mentioned in Section 5.7. For each website snippet, we attempt to locate the full sentence corresponding to the snippet by crawling the website’s content. However, the content of many websites could not be crawled. For such websites, we experiment with two strategies: (1) Discarding these websites during training and testing. (2) Using the incomplete snippets. The experimental results are shown in the table below, where “Raw” represents all snippets without completion, “Discard” represents strategy (1), and “Mixture” represents strategy (2). From the experimental results we can observe that directly discarding the websites leads to a significant performance loss, as discarding reduces the number of usable websites by approximately half, thereby limiting the performance of the website filter. Furthermore, as $\theta$ increases, the performance of strategy (2) becomes increasingly close to that of all snippets without completion (i.e., Raw), which validates that the completeness of the snippets has little impact on accuracy.
Table 5: Experiments of snippet completeness.
Baseline Variant $\theta$
10% 25% 40% 55% 70% 85% 100%
LLaVA-1.6 Raw 80.8 86.1 88.4 89.4 89.7 90.0 90.2
(Ours) Discard 76.6 80.2 80.7 82.1 81.9 81.3 81.6
Mixture 83.6 87.9 89.4 89.6 89.8 90.4 90.2
InternVL-1.5 Raw 84.6 89.2 91.1 92.9 92.4 92.7 92.9
(Ours) Discard 79.7 82.2 82.9 83.8 84.0 82.8 83.2
Mixture 86.0 89.5 92.2 92.3 92.2 92.5 92.9
@@ 6 Conclusion
In this work, we have presented SearchLVLMs, a plug-and-play framework to augment LVLMs in handling visual question answering about up-to-date knowledge. By introducing a hierarchical filtering model, the framework enables LVLMs to access up-to-date knowledge. A UDK-VQA dataset is further curated by scraping up-to-date news and generating news-related VQA samples. The dataset enables quantitatively evaluate the ability of LVLMs to respond to questions about up-to-date knowledge. Experimental results on UDK-VQA demonstrate that our framework can significantly boost the performance of LVLMs for answering questions requiring up-to-date knowledge.